The corners of a square lie on a circle of diameter D = 0.43 m. The side of the square has a length L. Find L.
Answer by lwsshak3(11628)

The corners of a square lie on a circle of diameter D = 0.43 m. The side of the square has a length L. Find L.
**
Draw a diagnol from one corner of the square to the opposite corner. This diagnol is equal to the diameter D=0.43 m, and is the hypotenuse of a right triangle with equal legs of L.
By Pythagroean Theorem:
.43^2=L^2+L^2=2L^2
L^2=.43^2/2=.0925
L=√.0925
L=.3041